# Diffusion models for multivariate subsurface generation and efficient probabilistic inversion

### v2

@@ 2.2.1 Corrected Diffusion Posterior Sampling (CDPS)
2:for $\sigma_{t}=0,...,\sigma_{max}$ do
3:  sample $\textbf{x}_{0}\sim q_{data}$
4:  $\textbf{x}_{t}\leftarrow\textbf{x}_{0}+\sigma_{\textbf{t}}\textbf{z},\textbf{z}\sim\mathcal{N}(0,I)$
5:  $\hat{\textbf{x}}_{0}^{(t)}\leftarrow S_{\theta}(\textbf{x}_{t},t)$
6:  $\sigma_{{\hat{x}_{0}}}^{(t)}\leftarrow\mathbb{E}\left[\sqrt{\sum_{i=1}^{N}(\hat{x}_{0,i}^{(t)}-x_{0,i})^{2}/N}\right]$
7:end for
8:sample $\textbf{x}_{T}\sim\mathcal{N}(0,I)$
9:for i = 0,…, N do
10:  $\sigma_{t}\leftarrow\sigma_{i}$
@@
11:  $\hat{\textbf{x}}_{0}^{(t)}\leftarrow S_{\theta}(\textbf{x}_{t},\sigma_{t})$
@@
12:  $\nabla_{x_{t}}\log p_{t}(\textbf{x}_{t})\leftarrow(S_{\theta}(\textbf{x}_{t},\sigma_{t})-\textbf{x}_{t})/\sigma_{\textbf{t}}^{2}$
13:  $\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\leftarrow\sigma_{\hat{\textbf{x}}_{0}}^{2^{(t)}}$
14:  if $\mathcal{F}$ is nonlinear then
15:   $\mathcal{J}_{\hat{\textbf{x}}_{0}}\leftarrow\mathcal{J}_{\mathcal{F}}(\hat{\textbf{x}}_{0}^{(t)})$
16:   $\tilde{\Sigma}_{\textbf{d}}\leftarrow\mathcal{J}_{\hat{\textbf{x}}_{0}^{(t)}}^{T}\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\mathcal{J}_{\hat{\textbf{x}}_{0}^{(t)}}+\Sigma_{\textbf{d}}$
17:  else
18:   $\tilde{\Sigma}_{\textbf{d}}\leftarrow\mathcal{F}^{T}\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\mathcal{F}+\Sigma_{\textbf{d}}$
19:  end if
@@
20:  $\nabla_{x_{t}}\log p_{t}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})\leftarrow-\nabla_{\textbf{x}_{t}}\left[(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})^{T}\tilde{\Sigma}_{\textbf{d}}^{-1^{(t)}}(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})\right]$
21:  $\nabla_{x_{t}}\log p_{t}(\hat{\textbf{x}}_{0}^{(t)}|\textbf{d})\leftarrow\nabla_{x_{t}}\log p_{t}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})+\nabla_{x_{t}}\log p_{t}(\textbf{x}_{t})$
22:  $x_{t-1}\leftarrow x_{t}+(\sigma_{t}-\sigma_{t-1})\nabla_{x_{t}}\log p_{t}(\hat{\textbf{x}}_{0}^{(t)}|\textbf{d})$
23:  if $\sigma_{\textbf{x}}(t)\not=0$ then
@@
24:   $\nabla_{x_{t}}\log p_{t}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})^{\prime}\leftarrow-\nabla_{\textbf{x}_{t}}\left[(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})^{T}\tilde{\Sigma}_{\textbf{d}}^{-1^{(t)}}(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})\right]$
25:   $\nabla_{x_{t}}\log p_{t}(\hat{\textbf{x}}_{0}^{(t)}|\textbf{d})^{\prime}=\nabla_{x_{t}}\log p_{t}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})+\nabla_{x_{t}}\log p_{t}(\textbf{x}_{t})$
26:   $x_{t-1}\leftarrow x_{t}+(\sigma_{t}-\sigma_{t-1})(\frac{1}{2}\nabla_{x_{t}}\log p_{t}(\hat{\textbf{x}}_{0}^{(t)}|\textbf{d})+\frac{1}{2}\nabla_{x_{t}}\log p_{t}(\hat{\textbf{x}}_{0}^{(t)}|\textbf{d})^{\prime})$
27:  end if
28:end for
29:Output: $\tilde{\textbf{x}}_{0}\sim p(\textbf{x}_{0}|\textbf{d})$
@@ 2.4 Performance assessment
Prior modeling performance is assessed by comparing the unconditional subsurface realizations against the TI and realizations generated with two benchmark models, a GAN and a VAE, proposed by Miele et al. (2024) for an analogous case study. The statistical distance between the target and modeled distributions (either prior or posterior) were quantified using the root-mean-square-error (RMSE) and the Kullback-Leibler divergence ($KL$) (Kullback and Leibler, 1951). The similarity of spatial features in the realizations was assessed using the structural similarity index (SSIM) (Wang et al., 2004) defined as
@@ 3.1 Unconditonal modeling
The joint distributions of facies and co-located $I_{P}$ obtained from the sampled realizations (Fig. 2b, d, and f) show significantly superior modeling ability of the DM (Fig. 2b). The facies classes are modeled with negligible errors (i.e., facies values different from 0 and 1); the corresponding $I_{P}$ marginal distributions (plotted along the y-axes) match accurately the prior means and standard deviations, fitting the data with a $KL$ divergence approximately five times lower than for GAN and VAE (Table 1). Further analysis of the networks unconditional modeling performances focusing on geostatistical metrics, are provided in Appendix A.
@@ 3.2.2 Nonlinear inversion results
The CDPS has superior prediction accuracy. For the whole study area, the generated ensemble captures the true $I_{P}$ much better than DPS, as shown by the lower values of logS in Table 4 and in the trace example in Figure 8h (x = 25 m). Moreover, each realization reproduces the target $I_{P}$-distribution with a low $KL$ divergence (Table 4) and presents fine-scale spatial variability patterns, which appear smoothed in DPS realizations (Fig. 8g).
@@ Appendix C CDPS implementation with DDPM and DDIM
2:for t = T,…, 1 do
@@
9:sample $\textbf{x}_{T}\sim\mathcal{N}(0,I)$
10:for t = T-1,…, 0 do
11:  $\epsilon_{\theta}^{(t)}\leftarrow S_{\theta}(\textbf{x}_{t},t)$
12:  $\hat{\textbf{x}}_{0}^{(t)}\leftarrow\left(\textbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\right)/\sqrt{\bar{\alpha}_{t}}$
13:  if $\mathcal{F}$ is nonlinear then
14:   $\mathcal{J}_{\hat{\textbf{x}}_{0}}\leftarrow\mathcal{J}_{\mathcal{F}}(\hat{\textbf{x}}_{0}^{(t)})$
15:   $\tilde{\Sigma}_{\textbf{d}}\leftarrow\mathcal{J}_{\hat{\textbf{x}}_{0}^{(t)}}^{T}\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\mathcal{J}_{\hat{\textbf{x}}_{0}^{(t)}}+\Sigma_{\textbf{d}}$
16:  else
17:   $\tilde{\Sigma}_{\textbf{d}}\leftarrow\mathcal{F}^{T}\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\mathcal{F}+\Sigma_{\textbf{d}}$
18:  end if
19:  $\nabla_{x_{t}}\log p_{t}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})\leftarrow-\nabla_{\textbf{x}_{t}}\left[(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})^{T}\tilde{\Sigma}_{\textbf{d}}^{-1}(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})\right]$
20:  $\hat{\epsilon}^{(t)}\leftarrow\epsilon_{\theta}^{(t)}+\sqrt{(1-\bar{\alpha}_{t})}\nabla_{x_{t}}\log p_{t}(\textbf{x}_{t})$
21:  $\textbf{z}\sim\mathcal{N}(0,I)$
22:  $x_{t-1}\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t-1}}\textbf{x}_{t}+\frac{\sqrt{\alpha_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t-1}}\left(\textbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\nabla_{x_{t}}\log p_{t}(\hat{\textbf{x}}_{0}^{(t)}|\textbf{d})\right)/\sqrt{\bar{\alpha}_{t}}+\sigma_{t}\textbf{z}$
23:end for
24:Output: $\tilde{\textbf{x}}_{0}\sim p(\textbf{x}_{0}|\textbf{d})$

### v3

@@ 2.2.1 Corrected Diffusion Posterior Sampling (CDPS)
2:
3:for $\sigma_{t}=0,...,\sigma_{max}$ do
4:  sample $\textbf{x}_{0}\sim q_{data}$ $\triangleright$ Evaluate denoiser uncertainty for the sampling noise levels
5:  $\textbf{x}_{t}\leftarrow\textbf{x}_{0}+\sigma_{\textbf{t}}\textbf{z},\textbf{z}\sim\mathcal{N}(0,I)$
6:  $\hat{\textbf{x}}_{0}^{(t)}\leftarrow S_{\theta}(\textbf{x}_{t},t)$
8:end for
9:
10:sample $\textbf{x}_{T}\sim\mathcal{N}(0,I)$
11:for i = 0,…, N do
12:  $\sigma_{t}\leftarrow\sigma_{i}$ $\triangleright$ Get $\sigma_{t}$ according to the sampling schedule (Eq. 22)
@@
13:  $\hat{\textbf{x}}_{0}^{(t)}\leftarrow S_{\theta}(\textbf{x}_{t},\sigma_{t})$ $\triangleright$ Evaluate the denoised image at noise level $\sigma_{t}$
@@
14:  $\nabla_{x_{t}}\log p_{t}(\textbf{x}_{t})\leftarrow(S_{\theta}(\textbf{x}_{t},\sigma_{t})-\textbf{x}_{t})/\sigma_{t}^{2}$ $\triangleright$ Evaluate denoising score at $\sigma_{t}$
15:
16:  $\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\leftarrow\sigma_{\hat{\textbf{x}}_{0}}^{2^{(t)}}$ $\triangleright$ Get denoiser uncertainty at $t$
17:  if $\mathcal{F}$ is nonlinear then $\triangleright$ Propagate denoiser uncertainty in the d domain
18:   $\mathcal{J}_{\hat{\textbf{x}}_{0}}\leftarrow\mathcal{J}_{\mathcal{F}}(\hat{\textbf{x}}_{0}^{(t)})$
19:   $\tilde{\Sigma}_{\textbf{d}}\leftarrow\mathcal{J}_{\hat{\textbf{x}}_{0}^{(t)}}^{T}\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\mathcal{J}_{\hat{\textbf{x}}_{0}^{(t)}}+\Sigma_{\textbf{d}}$
20:  else
21:   $\tilde{\Sigma}_{\textbf{d}}\leftarrow\mathcal{F}^{T}\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\mathcal{F}+\Sigma_{\textbf{d}}$
22:  end if
23:
@@
24:  $\nabla_{x_{t}}\log p_{t}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})\leftarrow-\nabla_{\textbf{x}_{t}}\left[(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})^{T}\tilde{\Sigma}_{\textbf{d}}^{-1^{(t)}}(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})\right]$ $\triangleright$ Evaluate likelihood score given $\textbf{x}_{0}^{(t)}$
25:  $\nabla_{x_{t}}\log p_{t}(\hat{\textbf{x}}_{0}^{(t)}|\textbf{d})\leftarrow\nabla_{x_{t}}\log p_{t}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})+\nabla_{x_{t}}\log p_{t}(\textbf{x}_{t})$ $\triangleright$ Evaluate posterior score
26:  $d\textbf{x}_{t}\leftarrow-\sigma_{t}\nabla_{x_{t}}\log p_{t}(\hat{\textbf{x}}_{0}^{(t)}|\textbf{d})$ $\triangleright$ Evaluate $d\textbf{x}/dt$ at $\sigma_{t}$
27:  $\textbf{x}_{t-1}\leftarrow x_{t}+(\sigma_{t-1}-\sigma_{t})d\textbf{x}_{t}$ $\triangleright$ Take Euler step from $\sigma_{t}$ to $\sigma_{t-1}$
28:
29:  if $\sigma_{t-1}\not=0$ then $\triangleright$ Apply $2^{\text{nd}}$ order correction
30:   $\hat{\textbf{x}}_{0}^{(t-1)}\leftarrow S_{\theta}(\textbf{x}_{t-1},\sigma_{t-1})$
31:   $\nabla_{x_{t-1}}\log p_{t-1}(\textbf{x}_{t-1})^{\prime}\leftarrow(S_{\theta}(\textbf{x}_{t-1},\sigma_{t-1})-\textbf{x}_{t-1})/\sigma_{{t-1}}^{2}$
32:   $\tilde{\Sigma}_{\textbf{d}}^{-1^{(t-1)}}\leftarrow\text{Repeat lines 13--19 for }{(t-1)}$
@@
33:   $\nabla_{x_{t-1}}\log p_{t-1}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t-1)})^{\prime}\leftarrow-\nabla_{\textbf{x}_{t-1}}\left[(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t-1)})-\textbf{d})^{T}\tilde{\Sigma}_{\textbf{d}}^{-1^{(t-1)}}(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t-1)})-\textbf{d})\right]$
34:   $\nabla_{x_{t-1}}\log p_{t-1}(\hat{\textbf{x}}_{0}^{({t-1})}|\textbf{d})^{\prime}=\nabla_{x_{t-1}}\log p_{t-1}(\textbf{d}|\hat{\textbf{x}}_{0}^{({t-1})})^{\prime}+\nabla_{x_{t-1}}\log p_{t-1}(\textbf{x}_{t-1})^{\prime}$
35:   $d\textbf{x}_{t}^{\prime}\leftarrow-\sigma_{t-1}\nabla_{x_{t-1}}\log p_{t-1}(\hat{\textbf{x}}_{0}^{({t-1})}|\textbf{d})^{\prime}$
36:   $x_{t-1}\leftarrow x_{t}+(\sigma_{t}-\sigma_{t-1})(\frac{1}{2}d\textbf{x}_{t}+\frac{1}{2}d\textbf{x}_{t}^{\prime})$
37:  end if
38:end for
39:Output: $\tilde{\textbf{x}}_{0}\sim p(\textbf{x}_{0}|\textbf{d})$
@@ 2.4 Performance assessment
Prior modeling performance is assessed by comparing the unconditional subsurface realizations against the TI and realizations generated with two benchmark models, a GAN and a VAE, proposed by Miele et al. (2024) for an analogous case study. The statistical distance between the target and modeled distributions (either prior or posterior) were quantified using the root-mean-square-error (RMSE) and the Kullback-Leibler divergence (KLKL) (Kullback and Leibler, 1951). The similarity of spatial features in the realizations was assessed using the structural similarity index (SSIM) (Wang et al., 2004) defined as
@@ 3.1 Unconditonal modeling
The joint distributions of facies and co-located $I_{P}$ obtained from the sampled realizations (Fig. 2b, d, and f) show significantly superior modeling ability of the DM (Fig. 2b). The facies classes are modeled with negligible errors (i.e., facies values different from 0 and 1); the corresponding $I_{P}$ marginal distributions (plotted along the y-axes) match accurately the prior means and standard deviations, fitting the data with a KLKL divergence approximately five times lower than for GAN and VAE (Table 1). Further analysis of the networks unconditional modeling performances focusing on geostatistical metrics, are provided in Appendix A.
@@ 3.2.2 Nonlinear inversion results
The CDPS has superior prediction accuracy. For the whole study area, the generated ensemble captures the true $I_{P}$ much better than DPS, as shown by the lower values of logS in Table 4 and in the trace example in Figure 8h (x = 25 m). Moreover, each realization reproduces the target $I_{P}$-distribution with a low KLKL divergence (Table 4) and presents fine-scale spatial variability patterns, which appear smoothed in DPS realizations (Fig. 8g).
@@ Appendix C CDPS implementation with DDPM and DDIM
2:for t = T,…, 1 do $\triangleright$ Evaluate the denoiser uncertainty at sampling noise levels
@@
9:
10:sample $\textbf{x}_{T}\sim\mathcal{N}(0,I)$
11:for t = T-1,…, 0 do
12:  $\epsilon_{\theta}^{(t)}\leftarrow S_{\theta}(\textbf{x}_{t},t)$ $\triangleright$ Evaluate the added noise at $t$
13:  $\hat{\textbf{x}}_{0}^{(t)}\leftarrow\left(\textbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\right)/\sqrt{\bar{\alpha}_{t}}$ $\triangleright$ Evaluate the denoised image at $t$
14:
15:  $\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\leftarrow\sigma_{\hat{\textbf{x}}_{0}}^{2^{(t)}}$ $\triangleright$ Get denoiser uncertainty at $t$
16:  if $\mathcal{F}$ is nonlinear then $\triangleright$ Propagate denoiser uncertainty in the d domain
17:   $\mathcal{J}_{\hat{\textbf{x}}_{0}}\leftarrow\mathcal{J}_{\mathcal{F}}(\hat{\textbf{x}}_{0}^{(t)})$
18:   $\tilde{\Sigma}_{\textbf{d}}\leftarrow\mathcal{J}_{\hat{\textbf{x}}_{0}^{(t)}}^{T}\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\mathcal{J}_{\hat{\textbf{x}}_{0}^{(t)}}+\Sigma_{\textbf{d}}$
19:  else
20:   $\tilde{\Sigma}_{\textbf{d}}\leftarrow\mathcal{F}^{T}\Sigma_{\hat{\textbf{x}}_{0}}^{(t)}\mathcal{F}+\Sigma_{\textbf{d}}$
21:  end if
22:
23:  $\nabla_{x_{t}}\log p_{t}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})\leftarrow-\nabla_{\textbf{x}_{t}}\left[(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})^{T}\tilde{\Sigma}_{\textbf{d}}^{-1}(\mathcal{F}(\hat{\textbf{x}}_{0}^{(t)})-\textbf{d})\right]$ $\triangleright$ Evaluate likelihood score at $t$
24:  $\hat{\epsilon}^{(t)}\leftarrow\epsilon_{\theta}^{(t)}+\sqrt{(1-\bar{\alpha}_{t})}\nabla_{x_{t}}\log p_{t}(\textbf{d}|\hat{\textbf{x}}_{0}^{(t)})$ $\triangleright$ Evaluate "posterior" noise at $t$
25:
26:  $\textbf{z}\sim\mathcal{N}(0,I)$
27:  $x_{t-1}\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t-1}}\textbf{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t-1}}\left(\textbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\nabla_{x_{t}}\log p_{t}(\hat{\textbf{x}}_{0}^{(t)}|\textbf{d})\right)/\sqrt{\bar{\alpha}_{t}}+\tilde{\sigma}_{t}\textbf{z}$ $\triangleright$ Take Euler step
28:end for
29:Output: $\tilde{\textbf{x}}_{0}\sim p(\textbf{x}_{0}|\textbf{d})$